The graph of an exponential function of the form y=a(b)^x passes through the points (1,2.8) and (2,3.92). What is the value of b?
Answer by stanbon(75887)

The graph of an exponential function of the form y=a(b)^x passes through the points (1,2.8) and (2,3.92). What is the value of b?
-------------
Substitute into y = a(b^x) to get two equations:
3.92=a*b^2
2.8 =a*b
------------
Divide the 2nd equation into the 1st to get:
b = 3.92/2.8
b = 1.4
============
Cheers,
Stan H.